# ReFRAME or Remain: Unsupervised Lexical Semantic Change Detection with Frame Semantics

### v1

@@ 5 Results
We evaluate our method on both subtasks of SemEval 2020 Task 1 on Unsupervised Lexical Semantic Change Detection. As described in Section 3, subtask 1 is a binary classification task evaluated by accuracy, while subtask 2 is a ranking task evaluated by Spearman’s rank correlation between the JSDs and the GOLD SCORES. Our primary focus is Subtask 2. For Subtask 1, we classify a word as changed if its JSD >= 0.5, else unchanged.
@@ 5.2 Subtask 1
The results for subtask 1 are presented in table 3. As the experiment set-up is fully unsupervised, different teams come up with different strategies to classify the changing status of the target lemmas. We based our binary classification score assignment on the 0.5 cutoff point of JSD. The threshold of 0.5 was chosen empirically, based on our prior experience using JSD on linguistic data, as a high-pass filter for semantic salience: it highlights substantial structural shifts while suppressing minor fluctuations that often arise from sparse distributions. We avoid any supervised finetuning for selecting the optimal JSD threshold to preserve the fully unsupervised spirit of the task and comparability with other unsupervised approaches. Both of our systems perform equally well, with an accuracy of 0.622, on par with different word embedding systems.
@@
6 Qualitative analysis
A key strength of our system is its interpretability. JSD decomposition reveals which distributional shifts drive the change, and users can inspect the corresponding contexts to locate the underlying semantic change. In this section, we qualitatively analyse the predictions of our better system (FTFE) to clarify what it captures.
@@ 6 Qualitative analysis
Upon qualitative analysis, semantic changes do indeed link to changes in frame distributions. The English noun prop develops a new theatre object meaning whereas in the earlier period the support meaning (either literal or figurative) was dominant. In the first period, prop were mostly linked to such frames as Possibility (e.g., the possibility of finding support) or Opinion (e.g, thinking about the pillar/bread-winner of the family or opposing the election of the prop/council of the city) or Building. In the second period, those frames decrease and there was an increase in Self motion frames, where prop acts as the mover (e.g., … props fly into scene) or looking for the prop is the purpose of the mover (e.g., he clambered … looking for perfect props) or Quantified_mass where prop was quantified by few, many or Bringing where prop is the object being brought. This is illustrated in figure 2.
@@
We only found one FP (quilt_nn). Our method detects large shifts in the frame distributions, yet it is ranked low in the gold Subtask 2 ranking. This is because of the large shift in the contexts (which do not correspond to any semantic change) in which quilt appears in. In the second period, quilt was the object of many activities, such as lean into the quilt. wring the quilt, lay the folded quilt, etc. This results in a large increase of such frames as Posture, Manipulation, Placing, Spatial contact, Causation, which drives the overall JSD up (see figure 4).
@@
There are two common problems we observe during the qualitative analysis. First, some frames are highly general and carry little specific semantic content. For instance, in the ball case, the rise of the Statement frame largely reflects the frequent occurrence of say in contexts that mention ball (whether referring to the social event or the object), rather than a substantive meaning shift. Second, using lemmatised corpora can occasionally disrupt grammatical agreement or sentence boundaries, which may confuse the parser and lead to incorrect frame assignments (this effect appears minor overall, as shown by the low JSD between the raw and lemmatised corpora’s frame inventories).
@@ 7 Conclusion
First, frame inventory and frame parsers availability are a hard constraint, making this method inapplicable to languages lacking robust frame parsers. Even for English, the frame parser is still not perfect. It cannot output the token offsets for frame elements, meaning we must use the lemmatised version of the corpora, which can affect (minorly) the performance of the approach.
Second, our qualitative analysis, while supporting the interpretability claim, simultaneously reveals one inherent limitation of distributional frame tracking. The method can overestimate change when large shifts in a target’s contexts do not correspond to genuine semantic change. Conversely, it can underestimate change when genuine semantic innovations are not reflected strong enough in frame distributional shifts.
@@
The data and all the codes used in this study is publicly available. No human data was used and no human participant was involved in the experiment.

### v2

@@ 5 Results
We evaluate our method on both subtasks of SemEval 2020 Task 1. As described in Section 3, subtask 1 is a binary classification task evaluated by accuracy, while subtask 2 is a ranking task evaluated by Spearman’s rank correlation between the JSDs and the GOLD SCORES. Our primary focus is Subtask 2. For Subtask 1, we classify a word as changed if its JSD >= 0.5, else unchanged.
@@ 5.2 Subtask 1
The results for subtask 1 are presented in table 3. As the experiment set-up is fully unsupervised, different teams come up with different strategies to classify the changing status of the target lemmas. We based our binary classification score assignment on the 0.5 cutoff point of JSD. The threshold of 0.5 was chosen empirically, based on our prior experience using JSD on linguistic data. It acts as a high-pass filter for semantic salience: it highlights substantial structural shifts while suppressing minor fluctuations that often arise from sparse distributions. We avoid any supervised finetuning for selecting the optimal JSD threshold to preserve the fully unsupervised spirit of the task and comparability with other unsupervised approaches. Both of our systems perform equally well, with an accuracy of 0.622, on par with different word embedding systems.
6 Qualitative analysis
A key strength of our system is its interpretability. JSD decomposition reveals which distributional shifts drive the change, and users can inspect the corresponding contexts to locate the underlying semantic change. In this section, we qualitatively analyse the predictions of our better system (FTFE) to investigate what has been captured.
@@ 6 Qualitative analysis
Upon qualitative analysis, semantic changes do indeed link to changes in frame distributions. The English noun prop develops a new theatre object meaning whereas in the earlier period the support meaning (either literal or figurative) was dominant. In the first period, prop were mostly linked to such frames as Possibility (e.g., the possibility of finding support) or Opinion (e.g, thinking about the pillar/bread-winner of the family, or opposing the election of the prop/council of the city) or Building. In the second period, those frames decrease and there was an increase in Self motion frames, where prop acts as the mover (e.g., … props fly into scene) or looking for the prop is the purpose of the mover (e.g., he clambered … looking for perfect props) or Quantified_mass where prop was quantified by few, many or Bringing where prop is the object being brought. This is illustrated in figure 2.
@@
We only found one FP (quilt_nn). Our method detects large shifts in its frame distributions, yet it is ranked low in the gold Subtask 2 ranking. This is because of the large shift in the contexts (which do not correspond to any semantic change) in which quilt appears in. In the second period, quilt was the object of many activities, such as lean into the quilt. wring the quilt, lay the folded quilt, etc. This results in a large increase of such frames as Posture, Manipulation, Placing, Spatial contact, Causation, which drives the overall JSD up (see figure 4).
@@
There are two common problems we observe during the qualitative analysis. First, some frames are highly general and carry little specific semantic content. For instance, in the ball case, the rise of the Statement frame largely reflects the frequent occurrences of say in contexts that mention ball (whether referring to the social event or the object), rather than a substantive meaning shift. Second, using lemmatised corpora can occasionally disrupt grammatical agreement or sentence boundaries, which may confuse the parser and lead to incorrect frame assignments (this effect appears minor overall, as shown by the low JSD between the raw and lemmatised corpora’s frame inventories).
@@ 7 Conclusion
First, frame inventory and frame parsers availability are a hard constraint, making this method inapplicable to languages lacking robust frame parsers. Even for English, the frame parser is still not perfect. It cannot output the token offsets for frame elements, meaning we must use the lemmatised version of the corpora, which can (minorly) affect the performance of the approach.
Second, our qualitative analysis, while supporting the interpretability claim, simultaneously reveals one inherent limitation of distributional frame tracking. The method can overestimate changes when large shifts in a target’s contexts do not correspond to genuine semantic change. Conversely, it can underestimate change when genuine semantic innovations are not sufficiently reflected in frame distributional shifts.
@@
The data and all codes used in this study are publicly available. No human data was used, and no human participant was involved in the experiment.